OpenStudy (anonymous):

what is the period and phase shift of y=-5cos(pi*x-3)+2

hartnn (hartnn):

for \(y = A \cos (\dfrac{2\pi}{T} x +\theta ) +b\) T is the period \(\theta \) is the phase shift

OpenStudy (anonymous):

so is the period 2 and the phase shift -3?

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

I was of the impression that it would be -3/pi.

OpenStudy (anonymous):

phase shift = c/b